# Gorenstein singularities with $\mathbb{G}_{m}$-action and moduli spaces of holomorphic differentials

### v1

@@ 8. Loci of Weierstrass points with fixed semigroups
$0\to\omega_{f}(-b_{i}P)\to\omega_{f}(-(b_{i}-1)P)\to\omega_{f}(-(b_{i}-1)P)\otimes\mathcal{O}_{P}\to 0$
@@
In general, $\mathcal{M}_{g,1}^{H}$ might be reducible. We denote by $\mathcal{M}_{g,1}^{H^{\circ}}$ an irreducible component of $\mathcal{M}_{g,1}^{H}$. We also denote by ¯$\overline{\mathcal{M}}_{g,1}^{H^{\circ}}$ its closure in the Deligne–Mumford compactification ¯$\overline{\mathcal{M}}_{g,1}$. We say that $\mathcal{M}_{g,1}^{H^{\circ}}$ is minimal, if every smooth pointed curve in the closure of $\mathcal{M}_{g,1}^{H}$ has the same semigroup as $H$, i.e., no further degeneration to more special gap sequences. For example, for $H=\{0,2,4,\ldots,2g-2,2g,2g+1,2g+2,\ldots\}$ in genus $g$, the locus $\mathcal{M}_{g,1}^{H}$ is minimal, parameterizing Weierstrass points in hyperelliptic curves.
@@
Let $G$ be a gap sequence of order $g$ such that $2g-1\in G$. Then an irreducible component $\mathcal{M}_{g,1}^{H^{\circ}}$ of $\mathcal{M}_{g,1}^{H}$ is an affine variety if and only if ¯$\overline{\mathcal{M}}_{g,1}^{H^{\circ}}\setminus\mathcal{M}_{g,1}^{H^{\circ}}$ is pure codimension $1$. In particular, if $2g-1\in G$ and if $\mathcal{M}_{g,1}^{H^{\circ}}$ is minimal, then $\mathcal{M}_{g,1}^{H}$ is affine.
@@
First, if the complement of $\mathcal{M}_{g,1}^{H^{\circ}}$ in ¯$\overline{\mathcal{M}}_{g,1}^{H}$ is not pure codimension $1$, then it follows from the general result [Goo69, Proposition 1] that $\mathcal{M}_{g,1}^{H}$ is not affine.
Conversely, suppose the complement of $\mathcal{M}_{g,1}^{H^{\circ}}$ in ¯$\overline{\mathcal{M}}_{g,1}^{H}$ is pure codimension $1$. Note that $\kappa_{1}+\psi$ is an ample divisor class on ¯$\overline{\mathcal{M}}_{g,1}$; see [Cor93, Theorem 2.2]. Additionally, by Theorem 8.3, $\kappa_{1}+\psi$ restricted to $\mathcal{M}_{g,1}^{H}$ is trivial. Therefore, the affinity of $\mathcal{M}_{g,1}^{H^{\circ}}$ follows since an ample divisor is supported on the divisorial boundary complement; see [Goo69, Proposition 3].
Moreover, note that ¯$\overline{\mathcal{M}}_{g,1}$ has a normal crossing boundary $\Delta$ of pure codimension $1$ parameterizing nodal curves. Therefore, the intersection of $\Delta$ with ¯$\overline{\mathcal{M}}_{g,1}^{H^{\circ}}$ remains to be pure codimension $1$. If $\mathcal{M}_{g,1}^{H^{\circ}}$ is minimal, then ¯$\overline{\mathcal{M}}_{g,1}^{H^{\circ}}\setminus\mathcal{M}_{g,1}^{H^{\circ}}=\Delta\cap\overline{\mathcal{M}}_{g,1}^{H^{\circ}}$ is pure codimension $1$, thus implying the last claim. ∎
We caution the reader that ¯$\overline{\mathcal{M}}_{g,1}^{H}$ might contain loci of smooth pointed curves $\mathcal{M}_{g,1}^{H^{\prime}}$, where $H^{\prime}$ is a more special semigroup than $H$. Therefore, affinity of $\mathcal{M}_{g,1}^{H}$ reduces to verifying the dimensions of such $\mathcal{M}_{g,1}^{H^{\prime}}$ and their containment relations.

### v2

@@ 8. Loci of Weierstrass points with fixed semigroups
$0\to\omega_{f}(-b_{i}P)\to\omega_{f}(-(b_{i}-1)P)\to\omega_{f}(-(b_{i}-1)P)\otimes\mathcal{O}_{P}\to 0.$
@@
In general, $\mathcal{M}_{g,1}^{H}$ might be reducible. We denote by $\mathcal{M}_{g,1}^{H^{\circ}}$ an irreducible component of $\mathcal{M}_{g,1}^{H}$. We also denote by $\overline{\mathcal{M}}_{g,1}^{H^{\circ}}$ its closure in the Deligne–Mumford compactification $\overline{\mathcal{M}}_{g,1}$. We say that $\mathcal{M}_{g,1}^{H^{\circ}}$ is minimal, if every smooth pointed curve in the closure of $\mathcal{M}_{g,1}^{H}$ has the same semigroup as $H$, i.e., no further degeneration to more special gap sequences. For example, for $H=\{0,2,4,\ldots,2g-2,2g,2g+1,2g+2,\ldots\}$ in genus $g$, the locus $\mathcal{M}_{g,1}^{H}$ is minimal, parameterizing Weierstrass points in hyperelliptic curves.
@@
Let $G$ be a gap sequence of order $g$ such that $2g-1\in G$. Then an irreducible component $\mathcal{M}_{g,1}^{H^{\circ}}$ of $\mathcal{M}_{g,1}^{H}$ is an affine variety if and only if $\overline{\mathcal{M}}_{g,1}^{H^{\circ}}\setminus\mathcal{M}_{g,1}^{H^{\circ}}$ is pure codimension $1$. In particular, if $2g-1\in G$ and if $\mathcal{M}_{g,1}^{H^{\circ}}$ is minimal, then $\mathcal{M}_{g,1}^{H}$ is affine.
@@
First, if the complement of $\mathcal{M}_{g,1}^{H^{\circ}}$ in $\overline{\mathcal{M}}_{g,1}^{H}$ is not pure codimension $1$, then it follows from the general result [Goo69, Proposition 1] that $\mathcal{M}_{g,1}^{H}$ is not affine.
Conversely, suppose the complement of $\mathcal{M}_{g,1}^{H^{\circ}}$ in $\overline{\mathcal{M}}_{g,1}^{H}$ is pure codimension $1$. Note that $\kappa_{1}+\psi$ is an ample divisor class on $\overline{\mathcal{M}}_{g,1}$; see [Cor93, Theorem 2.2]. Additionally, by Theorem 8.3, $\kappa_{1}+\psi$ restricted to $\mathcal{M}_{g,1}^{H}$ is trivial. Therefore, the affinity of $\mathcal{M}_{g,1}^{H^{\circ}}$ follows since an ample divisor is supported on the divisorial boundary complement; see [Goo69, Proposition 3].
Moreover, note that $\overline{\mathcal{M}}_{g,1}$ has a normal crossing boundary $\Delta$ of pure codimension $1$ parameterizing nodal curves. Therefore, the intersection of $\Delta$ with $\overline{\mathcal{M}}_{g,1}^{H^{\circ}}$ remains to be pure codimension $1$. If $\mathcal{M}_{g,1}^{H^{\circ}}$ is minimal, then $\overline{\mathcal{M}}_{g,1}^{H^{\circ}}\setminus\mathcal{M}_{g,1}^{H^{\circ}}=\Delta\cap\overline{\mathcal{M}}_{g,1}^{H^{\circ}}$ is pure codimension $1$, thus implying the last claim. ∎
We caution the reader that $\overline{\mathcal{M}}_{g,1}^{H}$ might contain loci of smooth pointed curves $\mathcal{M}_{g,1}^{H^{\prime}}$, where $H^{\prime}$ is a more special semigroup than $H$. Therefore, affinity of $\mathcal{M}_{g,1}^{H}$ reduces to verifying the dimensions of such $\mathcal{M}_{g,1}^{H^{\prime}}$ and their containment relations.